# Fast data inversion for high-dimensional dynamical systems from noisy measurements

### v2

@@ 4.2 Simulated experiments with known factor loadings
$\text{RMSE}_{s}$ $=\frac{1}{N}\sqrt{\frac{\sum_{t=1}^{n}\big{(}\mathbf{\hat{z}}_{s}(t)-\mathbf{z}_{s}(t)\big{)}^{T}\big{(}\mathbf{\hat{z}}_{s}(t)-\mathbf{z}_{s}(t)\big{)}}{k^{\prime}n}},$ (28)
where $\mathbf{\hat{z}}_{s}(t)=(\hat{z}_{s,1}(t),\dots,\hat{z}_{s,k^{\prime}})^{T}$ is the estimated slips at time $t$, computed by $\mathbf{\hat{z}}_{s}(t)=\mathbf{G}^{T}\mathbf{U}_{0}\mathbf{D}_{0}^{-2}\mathbf{\hat{z}}(t)$ with $\hat{\mathbf{z}}(t)$ being the estimation of latent factors.
@@
We compare FMOU with NIF in this section, as other previously compared methods, such as DMD, LY1 and LY5, estimate the factor loading matrix. We show the $\text{RMSE}_{m}$ and $\text{RMSE}_{s}$ by FMOU and NIF in Figure 6. As NIF does not involve the selection of the number of factors $d$, to eliminate the effect of selecting $d$, we assume $d$ is known for both FMOU and NIF. Across all configurations, the FMOU model consistently outperforms NIF in estimating the mean of the observations and the slip. Additionally, Figure S2 in supplementary materials compares the estimation of $d$ in Experiment 4 and it demonstrates that IC can correctly identify the number of factors with a moderately large number of time points. The improvement by FMOU for Experiment 4 is not surprising as data are generated by FMOU model. Next we use a misspecified model in Experiment 5 to illustrate the flexibility of FMOU estimation.
@@
$z_{s,h}(t)=z_{s,max}(1-E(\bm{\xi}_{h},t)^{2})\times\mathbbm{1}_{\{E(\bm{\xi}_{h},t)\leq 1\}},$ (29)
where $E(\bm{\xi}_{h},t)=(\xi_{h,1}-\xi_{0,1})^{2}/r_{0}^{2}+(\xi_{h,2}-\xi_{0,2})^{2}/r(t)^{2}$, $h=1,\ldots,k^{\prime},t=1,\ldots,n$. The slips have a peak value of $z_{s,max}=3$cm at the center and drop to zero outside the ellipse. We consider three configurations with known noise variances $\sigma_{0}^{2}\in\{0.01^{2},0.02^{2},0.05^{2}\}cm^{2}$.
In Experiment 5, both $\mathbf{U}_{0}$ and $\sigma_{0}^{2}$ are known to mimic the application of the real data. We use the VM method in (22) to estimate the number of latent factors. Table 4 records the numerical comparisons between FMOU and NIF for Experiment 5. We consider two variants of NIF: one with $d=k$ and the other one with an estimated $\hat{d}$ from IC. The FMOU achieves better accuracy in estimating both the mean of the observations and slips. Additionally, FMOU is considerably faster than the NIF model, as computing the likelihood function in FMOU only requires $\mathcal{O}(knd)$ operations. These orders do not consider SVD of Green’s function which only needs to be done once. Furthermore, as each step of EM algorithm has closed-form expressions, FMOU is also robust in estimating a large number of parameters.
@@
FMOU 105 $3.88\times 10^{-3}$ $2.25\times 10^{-1}$ $9.03$
NIF 105 $1.14\times 10^{-1}$ $6.41\times 10^{-1}$ $5.80\times 10^{2}$
NIF 200 $1.14\times 10^{-1}$ $6.41\times 10^{-1}$ $2.25\times 10^{3}$
@@
FMOU 67 $1.30\times 10^{-2}$ $3.37\times 10^{-1}$ $5.41$
NIF 67 $1.16\times 10^{-1}$ $6.57\times 10^{-1}$ $2.82\times 10^{2}$
NIF 200 $1.16\times 10^{-1}$ $6.57\times 10^{-1}$ $2.27\times 10^{3}$
@@
FMOU 30 $3.36\times 10^{-2}$ $4.82\times 10^{-1}$ $3.73$
NIF 30 $1.15\times 10^{-1}$ $6.67\times 10^{-1}$ $98.3$
NIF 200 $1.15\times 10^{-1}$ $6.60\times 10^{-1}$ $2.29\times 10^{3}$
Table 4: Results for Experiment 5. The standard deviation of the mean of the output and the slip are 0.19 and 0.85, respectively. The major cost in FMOU lies in estimating the number of latent factors. FMOU takes 8.13, 4.77 and 3.20 seconds for $\sigma_{0}^{2}=0.01^{2},0.05^{2},0.2^{2}$ to estimate the number of latent factors, $d$, respectively.
@@ 5.1 Data and methods
We employ GPS measurements to estimate slip rates within the Cascadia region, situated along the western edge of North America. The Cascadia region is characterized by the subduction of the Juan de Fuca plate beneath the North American plate. This geological interaction triggers slow slip events that last for weeks to months. The dataset used in this study is publicly available at the Plate Boundary Observatory, which contains observations from $\tilde{k}=100$ GPS stations over $n=88$ days in 2011 (Bartlow et al.,, 2011). Each observation represents the daily average geographical position recorded in three directions: East-West, North-South and Up-Down. We follow Bartlow et al., (2011) to use GPS measurements in the East-West and North-South directions, as the vertical displacement measurements contain little information for slip estimation. Figure 8(a) shows the cumulative displacements in the Cascadia region between June 3, 2011 to August 30, 2011, while the displacements in the East-West and North-South directions of six GPS stations are shown in Figure 8(b). As identified in DeMets et al., (2010), the observations encompass secular signals, such as annual and semiannual components, which are typically removed before the analysis (Bartlow et al.,, 2011). Details of preprocessing GPS measurements before modeling are provided in Section S6 in the supplement materials. Apart from the GPS measurements, we utilize determinations of tectonic tremor locations from the same region and period (Wech and Creager,, 2008) as an independent assessment of the location of fault slip at a given time.
@@
The Green’s functions used here are built on a triangular mesh consisting of $k^{\prime}=1978$ triangular patches (Fukushima et al.,, 2005) located on a fault plane beneath the ground and they are calculated by assuming triangular dislocations in a homogeneous and elastic half-space (Thomas,, 1993). Observations from 100 GPS stations in two directions yield a $200\times 1978$ Green’s function matrix $\mathbf{G}$. Furthermore, measurements are typically influenced by uncertainty in assigning the GPS-derived displacements in a North American Plate fixed reference frame, and local displacements of the GPS monuments. Thus it is common to include the frame motion or a time-dependent trend of each direction shared by each GPS station (Bartlow et al.,, 2011). Consequently, the augmented model can be written as $\mathbf{y}(t)={\mathbf{G}}^{aug}{\mathbf{z}}^{aug}_{s}(t)+\bm{\epsilon}(t)$, where ${\mathbf{G}}^{aug}=(\mathbf{G},\mathbf{I}_{k\times 2})$ with $\mathbf{I}_{k\times 2}:=(\mathbf{I}_{2},\dots,\mathbf{I}_{2})^{T}$, and ${\mathbf{z}}^{aug}_{s}(t)=[\mathbf{z}_{s}(t)^{T},f_{E}(t),f_{N}(t)]^{T}$ with $f_{E}(t)$ and $f_{N}(t)$ being the frame motion in the East-West and North-South directions, respectively, for $t=1,\ldots,n$. The Gaussian noise vector follows $\bm{\epsilon}(t)\sim N(0,\sigma_{0}^{2}\mathbf{I}_{k})$ with $\sigma_{0}^{2}$ being the observed variance. Following Bartlow et al., (2011), we estimate the magnitude of the fault slips projected onto a horizontal plane at 52 degrees clockwise from North along with fault direction (DeMets et al.,, 2010).
@@
${\mathbf{\hat{z}}}_{s,rate}(t)={\mathbf{\hat{z}}_{s}(t+1)-\mathbf{\hat{z}}}_{s}(t),$ (30)
for $t=1,\ldots,n-1$. Following Bartlow et al., (2011), we truncate the negative slip rates to zero, as it is deemed unlikely that the fault would slip backward in the prevailing tectonic stress state. For comparison, we include the NIF model (Segall and Matthews,, 1997) and the modified NIF model (Bartlow et al.,, 2011), which are summarized in Section S4.2 and Section S4.3 in the supplementary materials, respectively.
@@ 5.2 Results
To evaluate the model performance, we use the held-out tremor locations to validate the estimation, as tremor is well associated in both space and time with fault slip. The tremors are monitored continuously in the Cascadia zone (Wech and Creager,, 2008; Wech,, 2010), available from the Pacific Northwest Seismic Network (PNSN) catalog. We consider two metrics: the proportion of tremors identified by large slip rates, and the total areas containing large slip rates for evaluating the true positive rate and positive rate, respectively.
@@
The computational time of the three methods is given in panel (c) in Figure 9. Notably, the computational time of the FMOU is more than 300 times and 3000 times faster than the NIF model and the modified NIF model, respectively. The most computationally intensive step for the FMOU is estimating the number of factors, which takes 3.56 seconds. After selecting the number of latent processes, estimating all parameters and slip rates requires only 0.92 seconds. This dramatic acceleration in computation enables the use of massive datasets from large GPS networks, and provides new opportunities to jointly estimate the hazards across larger regions, which could otherwise be prohibitive due to the large computational expense.
@@
Figure 10 shows the seven-day average of the estimated slip rates of three periods by the FMOU, NIF and modified NIF models, where the solid dots represent the tremor epicenters. Estimation of slip rates of other periods is provided in Section S6 in supplementary materials. The estimates from FMOU align well with tremor dataset events, even though the GPS measurements and tremor datasets were collected independently from distinct sources. Compared to the modified NIF, the FMOU model provides better agreement between high-slip regions and tremor epicenters during August 10–16. In contrast, the modified NIF model detects a much larger area with relatively high slip rates, resulting in a higher false positive rate. Both FMOU and modified NIF reveal an area of high slip rates centered around $45^{\circ}$ and $123.5^{\circ}$ in early June, and another region with high slip rates centered around $47.5^{\circ}$ and $123^{\circ}$ in early August. However, the migration of the estimated high slip rates region is less apparent in the NIF model and its modified version. It’s worth mentioning that the finding from the FMOU model is consistent with previous work by Bartlow et al., (2011), which illustrates the physical mechanism of the coincidence between tremor centers and the regions with high slip rates.
@@ 6 Conclusion
In this paper, we proposed fast multivariate Ornstein-Uhlenbeck (FMOU) approach for estimating high-dimensional dynamical systems with noisy observations, We assumed a latent factor model with an orthogonal factor load matrix, where each latent process is modeled by an O-U process with distinct correlation and variance parameters estimated from the data. We derived a scalable EM algorithm in which each iteration contains closed-form expressions for parameter estimation and it does not require inverting large covariance matrix at each time in a Kalman filter. Extensive simulated studies and the real-world slip estimation example in the Cascadia region illustrate the high efficiency and scalability of the FMOU approach compared to other methods.
This study opens the door to various research problems in both methodology and applications. First, the O-U process is not mean-squared differentiable, so it may be of interest to extend the closed-form expressions for parameter estimation to differentiable Gaussian processes with Matérn covariance. Second, generalizing the fast algorithm to handle observations with irregular missing values is another interesting topic. Third, the FMOU implicitly induces a vector autoregressive model with a symmetric transition matrix in (26) for the mean of the observations. The high computational scalability and closed-form estimation of parameters make it appealing for applications such as scalable estimators for Granger causality from noisy data. For the application of slip estimation, the significantly faster approach by FMOU enables integrating geodetic data from large GPS networks, seismic data, and satellite interferograms for geological hazard quantification.

### v3

@@ 4.2 Simulated experiments with known factor loadings
(30) $\text{RMSE}_{s}$ $=\frac{1}{N}\sqrt{\frac{\sum_{t=1}^{n}\big{(}\mathbf{\hat{z}}_{s}(t)-\mathbf{z}_{s}(t)\big{)}^{T}\big{(}\mathbf{\hat{z}}_{s}(t)-\mathbf{z}_{s}(t)\big{)}}{k^{\prime}n}},$
where $\mathbf{\hat{z}}_{s}(t)=(\hat{z}_{s,1}(t),\dots,\hat{z}_{s,k^{\prime}}(t))^{T}$ is the estimated slips at time $t$, computed by $\mathbf{\hat{z}}_{s}(t)=\mathbf{G}^{T}\mathbf{U}_{0}\mathbf{D}_{0}^{-2}\mathbf{\hat{z}}(t)$ with $\hat{\mathbf{z}}(t)$ being the estimation of latent factors.
@@
We compare FMOU with NIF in this section, as other previously compared methods, such as DMD, LY1 and LY5, estimate the factor loading matrix. We show the $\text{RMSE}_{m}$ and $\text{RMSE}_{s}$ by FMOU and NIF in Figure 5. As NIF does not involve the selection of the number of factors $d$, to eliminate the effect of selecting $d$, we assume $d$ is known for both FMOU and NIF. Across all configurations, the FMOU model consistently outperforms NIF in estimating the mean of the observations and the slip. Additionally, Figure SM2 in supplementary materials compares the estimation of $d$ in Experiment 4 and it demonstrates that IC can correctly identify the number of factors with a moderately large number of time points. The improvement by FMOU for Experiment 4 is not surprising as data are generated by FMOU model. Next we use a misspecified model in Experiment 5 to illustrate the flexibility of FMOU estimation.
@@
(31) $z_{s,h}(t)=z_{s,max}(1-E(\bm{\xi}_{h},t)^{2})\times\mathbbm{1}_{\{E(\bm{\xi}_{h},t)\leq 1\}},$
where $E(\bm{\xi}_{h},t)=(\xi_{h,1}-\xi_{0,1})^{2}/r_{0}^{2}+(\xi_{h,2}-\xi_{0,2})^{2}/r(t)^{2}$, $h=1,\ldots,k^{\prime},t=1,\ldots,n$. The slips have a peak value of $z_{s,max}=3$cm at the center and drop to zero outside the ellipse. We consider three configurations with known noise variances $\sigma_{0}^{2}\in\{0.01^{2},0.05^{2},0.2^{2}\}cm^{2}$.
In Experiment 5, both $\mathbf{U}_{0}$ and $\sigma_{0}^{2}$ are known to mimic the application of the real data. We use the VM method in (22) to estimate the number of latent factors. Table 4 records the numerical comparisons between FMOU and NIF for Experiment 5. We consider two variants of NIF: one with $d=k$ and the other one with an estimated $\hat{d}$ from VM. The FMOU achieves better accuracy in estimating both the mean of the observations and slips. Additionally, FMOU is considerably faster than the NIF model, as computing the likelihood function in FMOU only requires $\mathcal{O}(knd)$ operations. These orders do not consider SVD of Green’s function which only needs to be done once. Furthermore, as each step of EM algorithm has closed-form expressions, FMOU is also robust in estimating a large number of parameters.
@@
FMOU 105 $0.0039$ $0.23$ $10$
NIF 105 $0.11$ $0.64$ $549$
NIF 200 $0.11$ $0.64$ $2271$
@@
FMOU 67 $0.013$ $0.34$ $6.0$
NIF 67 $0.12$ $0.66$ $251$
NIF 200 $0.12$ $0.66$ $2247$
@@
FMOU 30 $0.034$ $0.48$ $3.6$
NIF 30 $0.12$ $0.67$ $100$
NIF 200 $0.12$ $0.66$ $2243$
Table 4: Results for Experiment 5. The standard deviation of the mean of the output and the slip are 0.19 and 0.85, respectively. The major cost in FMOU lies in estimating the number of latent factors. FMOU takes 9.2, 5.3 and 3.2 seconds for $\sigma_{0}^{2}=0.01^{2},0.05^{2},0.2^{2}$ to estimate the number of latent factors, $d$, respectively.
@@ 5.1 Data and methods
We employ GPS measurements to estimate slip rates within the Cascadia region, situated along the western edge of North America. The Cascadia region is characterized by the subduction of the Juan de Fuca plate beneath the North American plate. This geological interaction triggers slow slip events that last for weeks to months. The dataset used in this study is publicly available at the Plate Boundary Observatory, which contains observations from $\tilde{k}=100$ GPS stations over $n=88$ days in 2011 [4]. Each observation represents the daily average geographical position recorded in three directions: East-West, North-South and Up-Down. We follow [4] to use GPS measurements in the East-West and North-South directions, as the vertical displacement measurements contain little information for slip estimation. Figure 7(a) shows the cumulative displacements in the Cascadia region between June 3, 2011 to August 30, 2011, while the displacements in the East-West and North-South directions of six GPS stations are shown in Figure 7(b). As identified in [10], the observations encompass secular signals, such as annual and semiannual components, which are typically removed before the analysis [4]. Details of preprocessing GPS measurements before modeling are provided in Section SM6 in the supplement materials. Apart from the GPS measurements, we utilize determinations of tectonic tremor locations from the same region and period [41] as an independent assessment of the location of geologic slip at a given time.
@@
The Green’s functions used here are built on a triangular mesh consisting of $k^{\prime}=1978$ triangular patches [11] located on a fault plane beneath the ground and they are calculated by assuming triangular dislocations in a homogeneous and elastic half-space [37]. Observations from 100 GPS stations in two directions yield a $200\times 1978$ Green’s function matrix $\mathbf{G}$. Furthermore, measurements are typically influenced by uncertainty in assigning the GPS-derived displacements in a North American Plate fixed reference frame, and local displacements of the GPS monuments. Thus it is common to include the frame motion or a time-dependent trend of each direction shared by each GPS station [4]. Consequently, the augmented model can be written as $\mathbf{y}(t)={\mathbf{G}}^{aug}{\mathbf{z}}^{aug}_{s}(t)+\bm{\epsilon}(t)$, where ${\mathbf{G}}^{aug}=(\mathbf{G},\mathbf{I}_{k\times 2})$ with $\mathbf{I}_{k\times 2}:=(\mathbf{I}_{2},\dots,\mathbf{I}_{2})^{T}$, and ${\mathbf{z}}^{aug}_{s}(t)=[\mathbf{z}_{s}(t)^{T},f_{E}(t),f_{N}(t)]^{T}$ with $f_{E}(t)$ and $f_{N}(t)$ being the frame motion in the East-West and North-South directions, respectively, for $t=1,\ldots,n$. The Gaussian noise vector follows $\bm{\epsilon}(t)\sim N(0,\sigma_{0}^{2}\mathbf{I}_{k})$ with $\sigma_{0}^{2}$ being the observed variance. Following [4], we estimate the magnitude of the geologic slips projected onto a horizontal plane at 52 degrees clockwise from North along with fault direction [10].
@@
(32) ${\mathbf{\hat{z}}}_{s,rate}(t)={\mathbf{\hat{z}}_{s}(t+1)-\mathbf{\hat{z}}}_{s}(t),$
for $t=1,\ldots,n-1$. Following [4], we truncate the negative slip rates to zero, as it is deemed unlikely that the fault would slip backward in the prevailing tectonic stress state. For comparison, we include the NIF model [33] and the modified NIF model [4], which are summarized in Section SM4.2 and Section SM4.3 in the supplementary materials, respectively. We do not include EM-VAR(1) in the real example as it requires inverting a $k^{\prime}\times k^{\prime}$ matrix for each time in the RTS smoother in the EM algorithm, which is too costly to compute.
@@ 5.2 Results
To evaluate the model performance, we use the held-out tremor locations to validate the estimation, as tremor is well associated in both space and time with geologic slip. The tremors are monitored continuously in the Cascadia zone [41, 40], available from the Pacific Northwest Seismic Network catalog. We consider two metrics: the proportion of tremors identified by large slip rates with overlapping grids (the true positive rate), and the total grids containing large slip rates (the positive rate).
@@
The computational time of the three methods is given in panel (c) in Figure 8. Notably, the computational time of the FMOU is more than 528 times and 1631 times faster than the NIF model and the modified NIF model, respectively. The most computationally intensive step for the FMOU is estimating the number of factors, which takes 3.03 seconds. After selecting the number of latent processes, estimating all parameters and slip rates requires only 0.54 seconds. This dramatic acceleration in computation enables the use of massive datasets from large GPS networks, and provides new opportunities to jointly estimate the hazards across larger regions, which could otherwise be prohibitive due to the large computational expense.
@@
Figure 9 shows the seven-day average of the estimated slip rates of three periods by the FMOU, NIF and modified NIF models, where the solid dots represent the tremor epicenters. Estimation of slip rates of other periods is provided in Section SM6 in supplementary materials. The estimates from FMOU align well with tremor dataset events, even though the GPS measurements and tremor datasets were collected independently from distinct sources. Compared to the modified NIF, the FMOU model provides better agreement between high-slip regions and tremor epicenters during August 10–16. In contrast, the modified NIF model detects a much larger area with relatively high slip rates, particularly in the southeast region during August 10–16, resulting in a higher false positive rate. Both FMOU and modified NIF reveal an area of high slip rates centered around $45^{\circ}$ and $123.5^{\circ}$ in early June, and another region with high slip rates centered around $47.5^{\circ}$ and $123^{\circ}$ in early August. However, the migration of the estimated high slip rates region is less apparent in the NIF model and its modified version. The finding by the FMOU model is consistent with the previous results by [4], which illustrates the physical mechanism of the coincidence between tremor centers and the regions with high slip rates.
@@ 6 Conclusion
In this paper, we proposed fast multivariate Ornstein-Uhlenbeck (FMOU) approach for estimating high-dimensional dynamical systems with noisy observations. We assumed a latent factor model with an orthogonal factor load matrix, where each latent process is modeled by an OU process with distinct correlation and variance parameters estimated from the data. We derived a scalable EM algorithm in which each iteration contains closed-form expressions for parameter estimation, and it does not require inverting the covariance matrix at each time in a Kalman filter as required in the conventional EM algorithm for the vector autoregressive model. Extensive numerical results illustrate the high efficiency and scalability of the FMOU approach compared to other methods.
This study opens the door to various research problems in both methodology and applications. First, the OU process is not mean-squared differentiable, so it is of interest to extend the closed-form expressions for parameter estimation to differentiable Gaussian processes with Matérn covariance. Second, generalizing the fast algorithm to handle observations with irregular missing values is another interesting topic. Third, the FMOU implicitly induces a vector autoregressive model with a symmetric transition matrix in (26) for the mean of the observations. The high computational scalability and closed-form estimation of parameters make it appealing for applications such as scalable estimators for Granger causality from noisy data. Finally, for the application of slip estimation, the significantly faster approach by FMOU enables integrating geodetic data from large GPS networks, seismic data, and satellite interferograms for geological hazard quantification.